Investment A has a beta of 1.7 and an expected rate of return of 15.7%. Investment B has a beta of 0.8 and an expected rate of return of 10.2%. What is the equity premium (market risk premium)?
Answer #1

SEE THE IMAGE. ANY DOUBTS, FEEL FREE TO ASK. THUMBS UP PLEASE

Home nert Page Layout Formulas Data Review View dd-Ins s Cut aCopy Σ AutoSum Calibri Wrap Text General ,_a. ars-函Merge & Center, $, % , 弼,8 C Paste B l u. Conditional Format CeInsert Delete Format Formatting, as Table w styles. ▼ ㆆ ▼ Sort &Find & 2 ClearFe Select Edting Format Painter Clipboard E0154 EB Font Alignment Number Styles Cells EC ED ЕЕ EF EG EH El EJ EK EL EM EN EO EP 150 151 152 153 154 155 156 157 158 159 160 161 162 163 164 165 166 167 168 4 KE CAPM UTILITY, SHARPE, beta bond Rf+beta(Rm-Rf) Rm-Rf RISK PREMIUM 15.7%-Af + 1.7(Rm-Rf) 10.2%-Af + 0.8( Rm-Rf) INVESTMENT A EQUATION 1 INVESTMENT B EQUATION 2 DOING (EQUATION 1 - EQUATION 2), WE WILL GET 15.7%-10.2% = 1.7(Rm-Rf) -0.8(Rm-Rf) 5.5%-0.9%( Rm-Rf) 6.1196-Rm-Rf MARKET RISK PREMIUM = Rm-Rf = 6.11% GM future INDEX INTL CAP BUDLEASING PV, FV, ANNUITYDIR cleanYIELD bond tru WACC RESI ex d 13096 06:34 14-01-2019